# Tokenizer Choice For LLM Training: Negligible or Crucial?

### v1

@@ 5.2.1 Impact of the Tokenizer Library
Table 2 demonstrates that BPE-SP, on average, outperforms BPE-HF across all tasks. In the multilingual setting, BPE-SP outperforms BPE-HF for all languages except for Spanish. The performance differences might be attributed to the differences in implementation details of the tokenisers’ pre-and postprocessing, which could affect the vocabulary creation (see Section 4.2) and, consequently, the downstream performance.
@@ 5.2.2 Impact of the Tokenizer Algorithm
Table 2 shows that for English tokenizers and models, BPE works better compared to Unigram. For multilingual models and tokenizers, we can see that for Spanish, German and French unigram can result in an improvement in performance. German and the romance languages are generally inflected languages, while English, although primarily analytic, does have some agglutinative features. Recent work showed Unigram tokenizers are better able to find morphologically accurate segmentations Bostrom and Durrett (2020); Park et al. (2021) which might help these languages.
@@ 5.2.3 Impact of the Tokenizer Vocabulary
Analyzing the impact of the vocabulary size revealed that in the monolingual English setting, the smaller/medium-sized vocabulary performs better (Table 3) whereas in the multilingual setting, larger vocabulary sizes result in better downstream performance(Table 3). This is in line with the vocabulary size of current English LLMs such as LLama.
@@ 5.3 Computational Costs
Even with an identical vocabulary size, two trained tokenizers most likely split a text passage in sequences of different lengths, resulting in different computation costs when processed by a model.
Our analysis revealed the following two insights. Firstly, increasing the vocabulary size decreases computational costs (Fig. 3). This implies that the reduction in the overall number of tokens represented by lower fertility (Fig. 1) compensates for the additional computational costs that a larger vocabulary size introduces in the model. In contrast, during inference, the computational costs increase with increasing vocabulary size and at least a vocabulary size of 50K, i.e., solely the forward pass without the backward pass does not compensate for the additional computational costs introduced by a larger vocabulary size in the model.
Secondly, the computational training costs for multilingual documents are significantly lower for multilingual tokenizers than monolingual English tokenizers for training (Fig. 2(a)) and for inference (in Appendix Fig. 8). In extreme cases, up to 68% additional training costs (see EN-UNI-SP-100352 vs. Multi-UNI-SP-50176 in Appendix Table 10 and Fig. 2(c)) as it is the case for training. At the same time, the difference in computational costs between multilingual and monolingual English tokenizers for processing English documents is much lower (Fig. 2(b)). A complete overview of the computational costs can be found in Table 10.
@@ 6 Correlation Between Intrinsic And Extrinsic Tokenizer Performance
Figure 5: Correlation of fertility/parity scores and downstream task performance for all five languages. We evaluated English models on English tasks, whereas multilingual models are evaluated across all four non-English tasks.
This section investigates a possible predictive relationship of intrinsic tokenizer metrics (fertilty and parity) to the extrinsic model downstream performance.
As highlighted in the correlation heatmaps in Fig. 5, we find that there is no distinct correlation across all tasks and languages, demanding a more granular analysis. While for non-English tasks, we mainly observe a correlation between low fertility and higher downstream performance, the non-English tasks yield seemingly random positive and negative correlations.
Taking the fertility trends with varying vocabulary sizes (see Fig. 1) into consideration, we hypothesize that fertility only correlates with downstream performance in certain language-specific vocabulary size limits. For the English language, the tokenizers already provide low, close-to-convergence fertility scores for vocabulary sizes of 32k tokens. While additional tokens yield only minute fertility improvements, we presume that they do not capture morphological segmentations and, thus, can harm downstream performance and significantly increase the computation costs (see Section 5.3) in the end.
@@
In conclusion, intrinsic tokenizer metrics such as fertility and parity need to be taken with a grain of salt and supposedly are only predictive of downstream model performance in certain bounds. Furthermore, the analysis showed that low fertility scores can be regarded as a necessary criterion but not as a sufficient one.
7 Conclusion
This work represents a fundamental step to a better understanding of the impact of the tokenizer on the models’ downstream performance. We have shown that training tokenizers with a balanced share across languages achieve comparable low fertility and parity scores across all languages, which has important implications. Lower fertility results in less computational costs and enables the model to learn long-range dependencies in limited context windows.
Furthermore, we highlight that the tokenizer choice can significantly impact the model’s downstream performance. We could show that it is beneficial to employ Unigram tokenizers with a large vocabulary size to train multilingual models. This is contrary to the current monolingual setting where BPE-based tokenizers are used together with a smaller/medium-sized vocabulary size. This also implies that insights obtained for monolingual settings are not transferable to multilingual settings and, therefore, require dedicated studies.
Finally, we could show that there is no clear correlation between intrinsic and extrinsic tokenizer performance, but the correlation is rather task-specific.
@@ Acknowledgements
This work was funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D) as well as by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine, LAMARR22B. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) as well as the Center for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for providing its facilities for high throughput calculations.
@@ Appendix A Corpora
Hyper-Parameter Value(s)
model_type Unigram |||| BPE
vocab_size 33k ||||50k |||| 82k |||| 100k
character_coverage 0.9999
split_by_number True
allow_whitespace_only_pieces True
add_dummy_prefix True
user_defined_symbols_1 <s>, </s>,<pad>,<eod>
user_defined_symbols_2 …, <placeholder_255>
byte_fallback True
max_sentence_length 4192
normalization_rule_name NFKC
train_extremely_large_corpus True
remove_extra_whitespaces False
split_by_whitespace True
Hyper-Parameter Value(s)
model_type BPE
vocab_size 33k ||||50k |||| 82k |||| 100k
limit_alphabet 512
use_nfkc_normalizer True
use_lowercase_normalizer False
use_strip_accents_normalizer True
pre_tokenizer ByteLevel, Digits
@@ Appendix C LLM Architecture and Hyperparameters
Regarding the training architecture of our 2.7B parameter models, we followed closely the architecture of GPT-3 Brown et al. (2020a). An overview of the used architecture details and hyperparameters is given in Table 8.
parameter value
@@ Appendix D Intrinsic Tokenizer Evaluation
Appendix E Extrinsic Tokenizer Evaluation
Accuracy
Model EN MULTI
BPE-HF-100 45.60 38.50
BPE-HF-32 45.34 38.12
BPE-HF-50 45.85 38.15
BPE-HF-82 44.97 38.67
BPE-SP-100 46.06 38.92
BPE-SP-32 47.06 38.01
BPE-SP-50 46.85 38.41
BPE-SP-82 45.25 38.64
UNI-SP-100 45.97 39.34
UNI-SP-32 46.62 38.18
UNI-SP-50 45.81 38.66
UNI-SP-82 45.65 39.21
Table 9: Average accuracy of monolingual and multilingual tokenizers across all downstream tasks
(a) Monolingual tokenizers
(b) Multilingual tokenizers
Figure 7: Accuracy distribution for mono-/multilingual tokenizers across mono-/multilingual tasks. The "+" represents the average performance across all tasks.
(a) Non-English documents.
(b) English documents.
(c) German documents.
Figure 8: Average compute (GFLOPs) required to process a single word within (a) multilingual, (b) English, and (c) German documents within an inference pass.
E.1 Performance Gains on Task-Level
In table Table 11, we show how much influence the choice of a tokenizer has on each downstream task. Therefore, we compare the results of the best- and worst-performing models for each task in our experiments. Performance increasements reach from +5,3% up to 380,9%, again showing the importance of choosing the right tokenizer on the downstream performance.
@@ E.1 Performance Gains on Task-Level
EN-BPE-HF-33 3,8 2,32 4,52
EN-BPE-HF-50 3,79 2,38 4,45
EN-BPE-HF-82 3,88 2,55 4,51
EN-BPE-HF-100 3,96 2,67 4,58
EN-BPE-SP-33 3,86 2,37 4,66
EN-BPE-SP-50 3,89 2,42 4,68
EN-BPE-SP-82 4,02 2,59 4,78
EN-BPE-SP-100 4,11 2,71 4,84
EN-UNI-SP-32 4,01 2,36 4,73
EN-UNI-SP-50 4,02 2,42 4,75
EN-UNI-SP-82 4,12 2,59 4,83
EN-UNI-SP-100 4,21 2,71 4,88
MULTI-BPE-HF-33 2,71 2,46 3,04
MULTI-BPE-HF-50 2,7 2,5 3,01
MULTI-BPE-HF-82 2,8 2,65 3,09
MULTI-BPE-HF-100 2,88 2,76 3,17
MULTI-BPE-SP-33 2,68 2,55 2,99
MULTI-BPE-SP-50 2,67 2,57 2,95
MULTI-BPE-SP-82 2,76 2,72 3,03
MULTI-BPE-SP-100 2,85 2,82 3,1
MULTI-UNI-SP-33 2,68 2,55 2,94
MULTI-UNI-SP-50 2,66 2,58 2,91
MULTI-UNI-SP-82 2,76 2,73 2,99
MULTI-UNI-SP-100 2,84 2,83 3,07
Table 10: Computational costs per word (GFLOPs) for different tokenizers.
Task Metric Worst Best Difference
EN arc challenge acc 0,21 0,27 +28,4 %
arc easy acc 0,50 0,59 +18,6 %
boolq acc 0,56 0,61 +8,9 %
cb acc 0,30 0,54 +76,5 %
cb f1 0,18 0,36 +94,9 %
copa acc 0,67 0,73 +9,0 %
hellaswag acc 0,34 0,41 +19,7 %
lambada openai acc 0,47 0,56 +20,0 %
mathqa acc 0,22 0,24 +8,8 %
mnli acc 0,35 0,37 +7,7 %
mnli mismatched acc 0,35 0,38 +9,4 %
mrpc acc 0,54 0,69 +26,1 %
mrpc f1 0,66 0,81 +22,7 %
gnad10 acc 0,15 0,43 +187,7 %
gnad10 f1 0,07 0,35 +380,9 %
gnad10 recall 0,14 0,42 +207,6 %
pawsx es acc 0,48 0,56 +17,2 %
piqa acc 0,67 0,72 +7,2 %
pubmedqa acc 0,45 0,59 +32,1 %
race acc 0,31 0,34 +11,8 %
rte acc 0,53 0,57 +8,9 %
sst acc 0,51 0,75 +48,2 %
triviaqa acc 0,02 0,05 +171,3 %
wic acc 0,50 0,55 +11,4 %
winogrande acc 0,49 0,56 +12,4 %
wnli acc 0,42 0,54 +26,7 %
wsc acc 0,36 0,61 +70,3 %
MULTI xstory cloze es acc 0,49 0,60 +24,3 %
mgsm de acc 0,00 0,03 -
lambada de acc 0,11 0,33 +186,4 %
lambada es acc 0,17 0,32 +91,8 %
lambada fr acc 0,20 0,38 +90,5 %
lambada it acc 0,16 0,37 +132,3 %
wino de acc 0,50 0,61 +20,9 %
xcodah de acc 0,27 0,42 +52,1 %
xcodah es acc 0,28 0,43 +52,1 %
xcodah fr acc 0,27 0,39 +43,7 %
xcsqa de acc 0,19 0,30 +56,4 %
xcsqa es acc 0,19 0,29 +57,3 %
xcsqa fr acc 0,17 0,29 +69,2 %
xnli de acc 0,32 0,48 +51,0 %
xnli en acc 0,49 0,52 +5,3 %
xnli es acc 0,34 0,48 +41,0 %
xnli fr acc 0,37 0,49 +30,3 %
xstance de acc 0,49 0,52 +6,0 %
xstance de f1 0,34 0,47 +41,2 %
xstance de recall 0,49 0,52 +7,7 %
Table 11: Differences between our best and worst performing tokenizers per task.

### v4

@@ 5.2.1 Impact of the Tokenizer Library
Table 5 demonstrates that BPE-SP, on average, outperforms BPE-HF in the monolingual and multilingual setting across all languages. The performance differences might be attributed to the differences in implementation details of the tokenizers’ pre-and postprocessing, which could affect the vocabulary creation (see Section 4.2) and, consequently, the downstream performance.
@@ 5.2.2 Impact of the Tokenizer Algorithm
Furthermore, Table 5 shows that depending on the language, either the BPE or Unigram exhibits better performance. It is noteworthy that the Germanic languages German and English benefit from the BPE algorithm, whereas the Romanic languages French and Spanish benefited from Unigram. The experiments for Italian, a Romanic language as well, show a different pattern than the other two Romanic languages.
@@ 5.2.3 Impact of the Tokenizer Vocabulary
Analyzing the impact of the vocabulary size revealed that in the monolingual English setting, the smaller/medium-sized, i.e., a vocabulary size of 33k/50k performs better (Table 5) whereas in the multilingual setting, in all cases except for German, larger vocabulary sizes result in better downstream performance. Taking into account the results presented in Table 3 showing that in the monolingual English setting, the best-performing tokenizer on average across all tasks had a vocabulary size of 33k and that the best-performing multilingual tokenizer had a vocabulary size of 100k additionally supports the observation that for the monolingual English setting a small vocabulary size is beneficial and for the multilingual setting a large vocabulary size is required.
@@ 5.3 Computational Costs
Given a fixed model, the computational costs depend on the vocabulary size and the fertility of the tokenizer, as defined in Eq. 1.
While larger vocabulary sizes introduce additional computational costs, they might also result in lower fertility scores and, therefore, lower overall computational costs for processing a set of documents, as discussed in Section 4. However, our findings in Fig. 3 show that increasing the vocabulary size from 50k to larger vocabulary sizes increases the computational costs in all cases. This highlights that the potentially lower fertility of larger vocabulary sizes cannot compensate for the additional costs introduced by the larger vocabulary size.
Furthermore, we observe that the computational training costs for multilingual documents are significantly lower for multilingual tokenizers than for monolingual English tokenizers (Fig. 2(a)). In fact, Fig. 2(b) and Table 11 in the appendix demonstrate that the training costs can increase up to 68% (comparing Multi-UNI-SP-50 to EN-UNI-SP-100 for German documents) for a given dataset. Assuming that during training it is required to process a fixed set of documents (e.g., Wikipedia to learn specific facts) entirely and not only a given number of tokens, the choice of the tokenizer can significantly impact the computational costs for training on this corpus.
While we could observe large cost differences between multilingual and monolingual English tokenizers in the monolingual English setting, the difference in computational costs between multilingual and monolingual English tokenizers for processing English documents is marginal (Fig. 2(c)).
@@ 6 Correlation Between Intrinsic And Extrinsic Tokenizer Performance
Figure 4: Spearman correlation of fertility/parity scores and downstream task performance for all five languages. We evaluated monolingual models on English tasks (left), whereas our multilingual models are evaluated across all non-English tasks. Pearson and Kendall correlation metrics showed a very similar picture.
This section investigates a possible predictive relationship of intrinsic tokenizer metrics (fertility and parity) to the extrinsic model downstream performance.
As highlighted in the correlation heatmaps in Fig. 4, we find that there is no distinct correlation across all tasks and languages, demanding a more granular analysis. While for non-English tasks, we mainly observe a correlation between low fertility and higher downstream performance, the non-English tasks yield seemingly random positive and negative correlations. However, it should be noted that the number of multilingual tasks per language is much lower than for English and that for several multilingual tasks such as XSQA and LAMBADA, a similar correlation behaviour between the English tasks and their translated version can be observed.
Taking the fertility trends with varying vocabulary sizes (see Fig. 1) into consideration, we hypothesize that fertility only correlates with downstream performance in certain language-specific vocabulary size limits. For the English language, the tokenizers already provide low, close-to-convergence fertility scores for vocabulary sizes of 33k tokens. While additional tokens yield only minute fertility improvements, we presume that they do not capture morphological segmentations and, thus, can harm downstream performance and significantly increase the computation costs (see Section 5.3) in the end.
@@
In conclusion, intrinsic tokenizer metrics such as fertility and parity need to be taken with a grain of salt and supposedly are only predictive of downstream model performance in certain bounds. Low fertility scores might be regarded as a necessary criterion but not as a sufficient one.
7 Conclusion & Future Work
This work represents a fundamental step to a better understanding of the impact of the tokenizer on the models’ downstream performance. We have shown that training tokenizers with a balanced share across languages achieve comparable low fertility and parity scores across all languages, which has important implications. Higher fertility results in up to 68% more computational costs during training and prevents the model from learning long-range dependencies in limited context windows.
Furthermore, we highlight that the tokenizer choice can significantly impact the model’s downstream performance. We could show that the BPE algorithm applies well to mono- and multilingual settings. For English, we show that a vocabulary size of 33k is sufficient, whereas multilingual models based on our five considered languages require a up to three times larger vocabulary size. Moreover, we could show that the SentencePiece library outperforms the Huggingface tokenizer library.
Finally, we could demonstrate that there is no clear correlation between intrinsic and extrinsic tokenizer performance, but the correlation is rather task-specific. A small fertility value might be a necessary condition for good downstream performance but not a sufficient one.
In the future, we aim to investigate tokenizers for a larger set of languages, including very diverse languages, and investigate the impact of alternative tokenization approaches such as SAGE Yehezkel and Pinter (2023) that focus on context information during tokenizer training.
8 Limitations
Despite the extensiveness of our work, it faces the following limitations.
Firstly, we did not perform hyper-parameter optimizations for each tokenizer. This was a deliberate choice to avoid additional computational costs, considering that training all 26 models only once required $\approx 59.000$ GPU hours.
Secondly, we did not investigate the effect of different random seeds on the model performance for a given tokenizer due to the additional computational costs. However, our results lay the foundation for future works that can further investigate the robustness of selected experiments.
Third, we did not investigate whether the results obtained could be extrapolated to larger model sizes, which we leave to future works. However, our finding that the BPE-SP-33 tokenizer is the best-performing tokenizer for the monolingual setting and the fact that this tokenizer has been used for training state-of-the-art models up to 65B Touvron et al. might indicate that our results also transfer to larger model sizes.
Finally, we did not provide results for a few-show setting since the metric of interest in the context of this work was the zero-shot downstream performance. Because we wanted to investigate whether the tokenizer choice impacts the model’s downstream performance, we argue that restricting on one of the widely applied metrics, i.e., the zero-shot setting, is sufficient to answer this research question. One further advantage of focusing on the zero-shot scenario is that we do not introduce an additional variable represented by the choice of the few-shot examples. However, we encourage future works to investigate whether our results translate into the few-shot evaluation setting.
9 Ethical And Broader Impact
LLMs represent a disruptive technology that has received significant attention from the public and is widely used across societies speaking different languages. Therefore, ensuring a democratization of the technology across people of different languages will represent an important value. Our study highlights that neglecting multilingualism while training a tokenizer representing a core component required for training LLMs can cause severe disadvantages, such as increased training costs and decreased downstream performance, raising major ethical concerns. Furthermore, the increased training costs translate into an increased carbon footprint, which has an environmental impact. Our findings support an improved development and usage of this fundamental technology.
@@ Acknowledgements
This work was funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D) as well as by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine, LAMARR22B and by the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101135671 (TrustLLM) and 952215 (TAILOR). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) as well as the Center for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for providing its facilities for high throughput calculations.
@@ Appendix B Tokenizer
Hyper-Parameter Value(s)
model_type Unigram |||| BPE
vocab_size 33k ||||50k
82k |||| 100k
character_coverage 0.9999
split_by_number True
allow_whitespace_only True
add_dummy_prefix True
user_symbols <s>,</s>,<pad>,
<eod>, <ph_1>,
…, <ph_255>
byte_fallback True
max_sentence_length 4192
normalization_rule_name NFKC
train_large_corpus True
remove_extra_whitespaces False
split_by_whitespace True
Hyper-Parameter Value(s)
model_type BPE
vocab_size 33k |||| 50k
82k |||| 100k
limit_alphabet 512
nfkc_normalizer True
lowercase_normalizer False
strip_accents_normalizer True
pre_tokenizer ByteLevel, Digits
Table 9: Overview of the Huggingface options that we used for the training of our tokenizers.
@@ Appendix C LLM Architecture and Hyperparameters
Regarding the training architecture of our 2.6B parameter models, we followed closely the architecture of GPT-3 Brown et al. (2020a). An overview of the used architecture details and hyperparameters is given in Table 10.
Hyper-Parameter Value
@@
For training the models, we used a fork of Megatron-LMhttps://github.com/NVIDIA/Megatron-LM.
@@ Appendix D Intrinsic Tokenizer Evaluation
D.1 Computational Costs Per Word During Training
Table 11 shows the average computational training costs for processing a word during the forward and backward pass.
@@ D.1 Computational Costs Per Word During Training
GPT-2-50 3.87 2.58 4.59
EN BPE-HF-33 3.8 2.32 4.52
BPE-HF-50 3.79 2.38 4.45
BPE-HF-82 3.88 2.55 4.51
BPE-HF-100 3.96 2.67 4.58
BPE-SP-33 3.86 2.37 4.66
BPE-SP-50 3.89 2.42 4.68
BPE-SP-82 4.02 2.59 4.78
BPE-SP-100 4.11 2.71 4.84
UNI-SP-32 4.01 2.36 4.73
UNI-SP-50 4.02 2.42 4.75
UNI-SP-82 4.12 2.59 4.83
UNI-SP-100 4.21 2.71 4.88
MULTI BPE-HF-33 2.71 2.46 3.04
BPE-HF-50 2.7 2.5 3.01
BPE-HF-82 2.8 2.65 3.09
BPE-HF-100 2.88 2.76 3.17
BPE-SP-33 2.68 2.55 2.99
BPE-SP-50 2.67 2.57 2.95
BPE-SP-82 2.76 2.72 3.03
BPE-SP-100 2.85 2.82 3.1
UNI-SP-33 2.68 2.55 2.94
UNI-SP-50 2.66 2.58 2.91
UNI-SP-82 2.76 2.73 2.99
UNI-SP-100 2.84 2.83 3.07
Table 11: Computational training costs per word (GFLOPs) for different tokenizers.
Appendix E Infrastructure & Computational Costs
We trained each of our 26 2.6B parameter models on NVIDIA A100 GPUs, and the training of each model took up to 2304 GPU hours. Therefore, the total training costs amounted to $\approx 59.000$ GPU hours.